imagine that you perform this experiment as described in your lab manual in doing so your first temperature reading was 33°C and your second reading was 71°C your first volume was 34 mL what would be your calculated second volume?
Answer #1

Given, T,= 33°c = 273 +33=306K Tz= tic = 273+ 7 = 344k N2= ? Vi= 34 ml charles law, from : V2 = _V, XT₂ = 34 x 344 306 Tue =thank you vmuch....